Payback Period and Accounting Rate of Return: Equal Annual Operating Cash Flows with Disinvestment
Minn is considering an investment proposal with the following cash flows:

Initial investment-depreciable assets $75,000
Net cash inflows from operations (per year for 10 years) 15,000
Disinvestment-depreciable assets 8,000

For parts b. and c., round answers to three decimal places, if applicable.

a. Determine the payback period.
Answer

years

b. Determine the accounting rate of return on initial investment.
Answer

c. Determine the accounting rate of return on average investment.
Answer
Answer #1
a)
Payback Period
a Initial Investment 75000
b Annual Cash Flow 15000
c Payback Period (a/b) 5 Years
b) accounting rate of return on initial investment.
Annual Depreciation = (75000-8000) /10 years
=6700
Annual Net Income = $15000-6700 =$8300
ARR =Net Income/Initinal Investment
=8300/$75000
11.07%
c) accounting rate of return on average investment.
=Net Income / (initinal investment + salvage value)/2
=8300/(75000+8000)/2
=$8300/41500
20.00%